OpenStudy (anonymous):

A group of people were given a personality test to determine if they were Type A or Type B. The results are shown in the table below. Male Female Type A 55 75 Type B 48 22

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

Compare P(Female or Type B) with P(Female | Type B). (5 points) There is not enough information. P(Female or Type B) > P(Female | Type B) P(Female or Type B) = P(Female | Type B) P(Female or Type B) < P(Female | Type B)

OpenStudy (anonymous):

Not sure what the question is asking : /

OpenStudy (kropot72):

To solve this you first need to sum the row and the column totals.

OpenStudy (anonymous):

Okay so Type A total is 130, Type B total is 70. Female is 97, Male is 103.

OpenStudy (kropot72):

P(F) = 97/200 P(M) = 103/200 P(A) = 130/200 P(B) = 70/200 \[\large P(F \cup B)=P(F)+P(B)-P(F \cap B)=you\ can\ calculate\]

OpenStudy (anonymous):

And this will tell me the answer?

OpenStudy (kropot72):

Compare the result with: \[\large P(F|B)=\frac{22}{70}\]

OpenStudy (anonymous):

i'm really confused @kropot72

OpenStudy (anonymous):

@kropot72 how do i find P(F upside down U thing B)?

OpenStudy (kropot72):

\[\large P(F \cap B)=P(F) \times P(B)\]

OpenStudy (anonymous):

ahh okay, let me try and figure it out

OpenStudy (anonymous):

so 167/200 - 6790?

OpenStudy (kropot72):

No. Where did you get 6790 from?

OpenStudy (anonymous):

P(f) x P(b)

OpenStudy (kropot72):

\[\large P(F) \times P(B)=\frac{97}{200}\times\frac{70}{200}=\frac{97\times70}{200\times200}=?\]

OpenStudy (anonymous):

OH okay hold on

OpenStudy (anonymous):

so it'd be 167 / 200 - 0.16975?

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

Correct.

OpenStudy (anonymous):

Okay, and so how does this exactly relate to the answer? @kropot72

OpenStudy (anonymous):

@kropot72 !

OpenStudy (kropot72):

You need to compare the results of: P(female or type B) which is written in correct symbols as: \[\large P(F \cup B)\] and P(female | given type B) which is written in correct symbols as: \[\large P(F|B)\] So you need to finish the two calculations, and then choose the correct answer.

OpenStudy (anonymous):

what exactly does P(F|B) mean? and is the answer D?

OpenStudy (anonymous):

@

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

As I told you in my previous post: P(female | given type B) is written in correct symbols as: \[\large P(F|B)\] Why don't you know these symbols if you are studying stats at this level?

OpenStudy (anonymous):

It's a remedial course and I had to leave for a week to volunteer, so I'm back now and this specific type of question was always hard for me. :(

OpenStudy (anonymous):

I don't understand what the | is though. @kropot72

OpenStudy (kropot72):

It is a symbol for writing the conditional probability that B occurs if A has occurred in the following way: P(B|A) It can be read as 'the probability of B given A'.

OpenStudy (anonymous):

okay wait! i think i have it. @kropot72 it's C right?

OpenStudy (anonymous):

because p(f|b) is 48, and the other choice is more than 48. does that sounds right?

OpenStudy (kropot72):

The maximum possible value of probability is 1.0000. I already posted that: \[\large P(F|B)=\frac{22}{70}=you\ can\ calculate\]

OpenStudy (anonymous):

ok and so this is where i'm not too sure where to go?

OpenStudy (anonymous):

the other is P(F or B), so wouldn't that clearly be bigger? @kropot72

OpenStudy (anonymous):

i'm sorry if this is frusturating for you.

OpenStudy (kropot72):

\[\large (F \cup B)=0.835-0.17=0.665\] \[\large P(F|B)=0.314\] \[\large P(F \cup B) >P(F|B)\] Please post the exact wording of your choice of answer.

OpenStudy (anonymous):

There is not enough information. P(Female or Type B) > P(Female | Type B) P(Female or Type B) = P(Female | Type B) P(Female or Type B) < P(Female | Type B)

OpenStudy (anonymous):

Okay, so I think I was right? The answer is B??

OpenStudy (kropot72):

Yes, the second choice of answer is correct. "P(Female or Type B) > P(Female | Type B"

OpenStudy (anonymous):

Thank you so much!!! :)

OpenStudy (kropot72):

You're welcome :)